Circular No. 8052

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2002kh, 2002ki, and 2003N S. Casertano, Space Telescope Science Institute, reports, on behalf of

the GOODS Treasury Team (cf. IAUC 7981, 8012, and 8038, plus V. Laidler, with the support of ground-based images from P. Capak and L. Cowie) and the Hubble Higher-z Supernova Team (+ M. Hicken), the discovery of three new apparent supernovae in images of the region around the Hubble Deep Field, each detected by subtracting ACS images taken on 2002 Nov. 21–22 from those taken on 2003 Jan. 2–4. SN 2003N was not obviously visible in the Nov. images and was clearly visible in the Jan. images; SN 2002ki was substantially brighter in the Nov. images but remained visible in the January images; and SN 2002kh was not detected in the Jan. 2–4 images, but was clearly visible earlier. Each apparent supernova was detected independently in each of the four separate z-band (F850lp) exposures taken at the relevant epoch, as well as in the two i-band (F775W) and the two v-band (F606W) exposures. The offsets given below are from the centers of the nearest galaxies visible in the ACS images. The magnitude quoted for SN 2003N assumes no supernova light in the earlier image. Further z-band magnitudes: SN 2003N, 2002 Nov. 21.53 UT, [27.5; SN 2002ki, 2003 Jan. 4.61 UT, > 25.1 (uncertain due to unknown contribution of host-galaxy light); SN 2002kh, 2003 Jan. 4.31, [27.0.

SN  2002/03  α_{2000}  δ_{2000}  z  Offset
2002kh Nov. 22.31 12^{h}36^{m}16.^{s}78 +62^{o}14′37′′.7 25.1 1′′.2 W, 0′′.3 N
2002ki Nov. 22.69 12 37 28.35 +62 20 40.0 24.2 0′′.1 W, 0′′.0 N
2003N Jan. 2.83 12 37 09.06 +62 11 00.7 25.9 0′′.0 W, 0′′.2 N

NOVA IN M31 A nova in M31 (= NGC 224) has been reported independently by the

LOTOSS group (M. Papenkova, J. Graham, and W. Li, University of California at Berkeley; cf. IAUC 7906) and by a Padua group (M. Fiaschi, F. Di Mille, R. Cariolato, and A. Frassati; cf. IAUC 7709). The LOTOSS group reports a precise position of α = 0^{h}42^{m}15^{s}.75, δ = +41^{o}13′24′′.6 (equinox 2000.0), which is 323′′ west and 164′′ south of the center of M31. Fiaschi et al. report position end figures 15^{s}.77, 24′′.7. Available magnitudes (K = from unfiltered KAIT CCD images; H = Hα magnitudes from Fiaschi et al.): 2002 Nov. 10, [18 (H); 2003 Jan. 5 UT, [19.0 (K); 8.1, 16.5 (K); 11.73, 15.8 (H); 14.73, 16.4 (H).

2003 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green